how many years will it take for $490 to grow to 993.78 at 8 percent
Answer by Alan3354(69443)

how many years will it take for $490 to grow to 993.78 at 8 percent
---------------
Depends on the compounding period.
---
FV = PV*(1+r)^n
If it's compounded annually:
993.78 = 490*1.08^n (n in years)
1.08^n = 993.78/490
n*log(1.08) = log(993.78/490)
n = log(993.78/490)/log(1.08)
n =~ 9.1879 years